# Sparse Ellipsoidal Radial Basis Function Network for Point Cloud Surface Representation

### v2

@@ 3.4.3 Adaptive Basis Function Addition
We introduce an adaptive basis function addition mechanism that automatically adds suitable basis functions to areas with large errors in the process of optimization of loss function. Before adding new basis functions, two prerequisites must be satisfied: (1) the condition in Eq. (12) and Eq. (13) have been fulfilled, meaning the $\mathcal{L}_{1}$ is active; (2) the number of effective basis functions remains stable over a predefined number of optimization steps, formulated as:
@@
where $b_{f}^{(e)}$ indicates the number of effective basis functions at epoch $e$, and is computed as: $b_{f}^{(e)}=\sum_{j=1}^{M}\mathbb{I}\left(\left|w_{j}\right|\geq\tau_{d}\right)$, $\mathbb{I}(\cdot)$ is the indicator function, which returns 1 if the condition is true and 0 otherwise. The threshold $\tau_{d}$ is used to determine whether a basis function is effective. $k^{l_{1}}$ and $\tau^{l_{1}}$ represent the number of iterations used to determine the stability of the number of basis functions and the corresponding threshold, respectively. Once the condition of Eq. (19) is triggered, the basis function addition process is automatically performed. After adding the new basis functions, the $\mathcal{L}_{1}$ regularisation term is temporarily deactivated and the network switches to pure $\mathcal{L}_{2}$ optimisation to fully train the newly added basis functions. The $\mathcal{L}_{1}$ regularisation term is reactivated for sparse optimization when Eq. (12) and Eq. (13) are satisfied again. The process repeats as follows:
@@
    When the condition for $\mathcal{L}_{1}$ activation is satisfied (Eq. (12) and Eq. (13)), $\mathcal{L}_{1}$ regularization is reactivated for sparse optimization.
@@
The process alternates cyclically until the predefined number of training epochs $T^{e}$ is reached. Afterward, the optimisation focuses exclusively on the $\mathcal{L}_{2}$ loss term to further improve the precision of the SDF approximation, continuing until the maximum iteration count $M^{e}$ is reached. At this stage, no further basis function additions or deletions occur, and $\mathcal{L}_{1}$ regularization is no longer applied.
The detailed process for adding basis functions is given in Algorithm 2. First, calculate the error vector for each training point in $V^{ls}$, and select the subset indexed by $id_{all}$, which is computed using Eq. (10):
$\bm{A}^{a}=\mathbf{0},$
where $\mathcal{I}_{ext}$ denotes the indices of the extreme error points. The length of $\mathcal{I}_{ext}$ is $k^{a}$, which indicates the number of newly added basis functions. $\bm{C}^{a}\in\mathbb{R}^{k^{a}\times 3}$ represents the centers of the newly added basis functions. $\bm{W}^{a}\in\mathbb{R}^{k^{a}\times 1}$ represents the corresponding weights. $\bm{d}^{a}\in\mathbb{R}^{k^{a}}$ represents the axis lengths of the newly added basis functions. Since the axis lengths are initialized to be identical along all three directions, each scalar element in $\bm{d}^{a}$ is expanded into a $3\times 3$ diagonal matrix to construct $\bm{D}^{a}\in\mathbb{R}^{k^{a}\times 3\times 3}$. $\bm{A}^{a}\in\mathbb{R}^{k^{a}\times 3}$ denotes the rotation angles, which are initialized to zero. $\bar{d}_{i}$ refers to the $i$-th element of $\bm{\bar{d}}$, which is defined as the minimum distance from the center $\bm{c}^{a}(i)$ to the point cloud $P$, $\bar{d}_{i}=\min\left\|\bm{c}^{a}(i)-P\right\|,i=1,2,\ldots,k^{a}$. $\operatorname{sign}()$ is the sign function:
$\operatorname{sign}(x)=\begin{cases}-1,&x<0\\ 0,&x=0\\ 1,&x>0\end{cases}$
@@
1:Identify high-error points: $\mathcal{I}_{add}=\{i\mid|E_{i}|>\sqrt{\frac{\tau_{m}}{2}}\}$.
@@
11:     return ,,$\emptyset$,$\emptyset$,$\emptyset$,$\emptyset$ , ∅ , ∅ , ∅
@@
During SE-RBFNet optimization, some basis functions have coefficients that become so small that their contribution to the overall error can be ignored. We introduce a threshold $\tau_{d}$ and remove the basis functions whose $w$ coefficients are less than $\tau_{d}$ at the specified iteration interval. That is:
$\bm{\widetilde{W}}$ =$=\left\{w_{j}\in\bm{W}|\>|w_{j}|<\tau_{d}\right\},$ (24)
$\bm{\widetilde{C}}$ =$=\left\{\bm{c}_{j}\in\bm{C}|\>|w_{j}|<\tau_{d}\right\},$
$\bm{\widetilde{D}}$ =$=\left\{\bm{D}_{j}\in\bar{\bm{D}}|\>|w_{j}|<\tau_{d}\right\},$
$\bm{\widetilde{A}}$ =$=\left\{\bm{a}_{j}\in\bm{A}|\>|w_{j}|<\tau_{d}\right\}.$
Deleting these basis functions allows the network to automatically adjust its complexity, increasing sparsity and computational efficiency.
3.4.4 Parallel Computing Acceleration
Finally, as in pratice, the sizes of the input set $G$ contains a very large number of points, and the problem of optimization $\mathcal{L}$ is a high dimensional non-convex nonlinear problem. To reduce computational burden and increase convergence speed, we primarily employ the following techniques for optimization.
@@ 3.4.4 Parallel Computing Acceleration
$\bm{F}_{ij}=\begin{cases}\hat{\Phi}\left(\bm{v}_{i},\Theta_{j}\right),&\text{ if }\left\|\bm{c}_{j}-\bm{v}_{i}\right\|_{2}<\sqrt{\frac{-\ln(\varepsilon)}{\lambda_{j}}}\\ 0,&\text{ otherwise }\end{cases}$ (26)
@@
The detailed derivations of the gradients can be found in the A. As shown in the A, when the value of $\bm{F}$ at a certain location is close to zero, the corresponding gradient also approaches zero. Consequently, to improve the calculation efficiency, we take advantage of the nearest-neighbor strategy as described in Eq. (25), only the gradient values of the points near the center points of the basis functions are calculated. Simultaneously, due to the inherent independence of each ERBF in the network, we leverage CUDA to parallelize the computation of value of the loss function and the gradient with respect to each parameter in $\Theta$. This independence allows each ERBF to be computed independently and concurrently, leading to a significant reduction in computational time.
@@
Algorithm 3 SE-RBFNet Sparse Optimisation Algorithm
Input: A point cloud containing the surface $\hat{P}$, batch size $B_{s}=10000$, $M^{e}=2000$, $T^{e}=1600$, initial learning rate $lr=0.01$, maximum depth of the octree $l_{max}=10$, $l_{s}$ is set by default to the third-to-last layer, $\gamma=10^{-3}$, $\varepsilon=10^{-7}$, $\tau_{1}=0.9$, $\tau_{2}=1.1$, $\tau_{m}=0.0005$, $\tau_{d}=0.01$, $\tau^{l_{1}}=10$, $\tau^{l_{2}}=0.5$, $k^{l_{1}}=50$, $k^{l_{2}}=10$. 1:Sample point cloud points $P$ and construct an ocree to obtain grid points $G$; 2:Calculate the signed distance from grid point $G$ and normalise according to Eq. (6) to obtain $\hat{S}$; 3:According to the $l_{s}$, obtain the initial training points $V^{l_{s}}$ and corresponding real labels $T^{ls}$, and then calculate the parameters of the initial ERBF using Algorithm 1. 4:Initialize SE-RBFNet parameters: $\bm{\widetilde{\Theta}}\leftarrow\{\bm{C}\textrm{,}\bar{\bm{D}}\textrm{,}\bm{A},\bm{W}\}$. 5:Set $l_{1}\_{optim}$=False, $add\_point$=False, $e_{p}=1$; 6:while $e_{p}<M^{e}$ do 7:     $e_{p}\leftarrow e_{p}+1$, 8:     if $add\_point$ and $l_{s}<l_{max}$ then 9:         The grid points from the $(l_{s}+1)$-th layer are selected and appended to the training set, followed by updating: $V^{l_{s}}\leftarrow V^{l_{s}+1}$, $T^{ls}\leftarrow T^{ls+1}$ and $l_{s}\leftarrow l_{s}+1$. 10:         Set $add\_point$ = False 11:     end if 12:     Shuffle training points and corresponding labels 13:     for $i=0$ to $N$ with step $B_{s}$ do 14:         Sample batch $V^{l_{s}}_{b}$ and $T^{l_{s}}_{b}$ 15:         Use Eq. (5) to calculate the network forward result: $\bm{O}\leftarrow\text{SE-RBFNet}(V^{l_{s}}_{b})$ 16:         Calculate $id_{all}$ according to Eq. (10) 17:         Compute $\mathcal{L}_{2}$ and $\mathcal{L}_{1}$ loss using Eq. (9) 18:         Explicitly compute the gradient $\nabla{\mathcal{L}(\bm{\widetilde{\Theta}})}$ using the formulas in the A. 19:         if $l_{1}\_{optim}$=True and Eq. (13) is True then 20:              Use Eq. (14) to compute adaptive $\alpha,\beta$ 21:              Update loss: $\mathcal{L}\leftarrow\alpha\mathcal{L}_{2}+\beta\mathcal{L}_{1}$ 22:         else 23:              Update loss: $\mathcal{L}\leftarrow\mathcal{L}_{2}$ 24:         end if 25:         Update network parameters $\bm{\widetilde{\Theta}}$ via Adam [63]. 26:     end for 27:     if $e_{p}\>\%\>k^{l_{2}}=0$ and $l_{1}\_{optim}$=True and $e_{p}<T^{e}$ then 28:         Delete invalid basis functions using Eq. (24). 29:     end if 30:     if $l_{1}\_{optim}$=False and Eq. (12) is True then 31:         Update $l_{1}\_{optim}$=True. 32:     end if 33:     if $l_{1}\_{optim}$=True and Eq. (19) is True then 34:         Add new ERBF accroding Eq. 21. 35:         Update $l_{1}\_{optim}$=False. 36:     end if 37:     if $e_{p}=T^{e}$ then 38:         During the iterations from $T^{p}$ to $M^{p}$, steps 24 to 30 are skipped, and only $\mathcal{L}_{2}$ optimisation is performed. $l_{1}\_{optim}$=False, $lr=10^{-3}$ and adjusted using a cosine annealing schedule, with a minimum value of $10^{-5}$. 39:     end if 40:     if $e_{p}$ in $[400,800]$ then 41:         $add\_points=\text{True}$ 42:     end if 43:end while Output: Network parameters $\bm{\widetilde{\Theta}}$, predicted SDF $\bm{O}$, a surface via the marching cubes algorithm [23].
@@ 4 Experiment and analysis
In this section, to evaluate the accuracy, sparsity and efficiency of our method (SE-RBFNet), we compare it with SparseRBF[1], which similarly represents surfaces using sparse radial basis functions. It is important to clarify that SE-RBFNet does not directly perform surface reconstruction from point clouds. Instead, it approximates the SDF, which allows the implicit surface representation as a combination of a relatively small number of ellipsoids. The surface can then be naturally reconstructed from the simplified and approximated SDF. Therefore, we do not directly compare with related surface reconstruction methods. All experiments were performed on Ubuntu 18.04 with an Intel Core i9-7940X CPU @ 3.1GHz and an Nvidia GeForce RTX 2080 GPU.
@@ 4.1 Datasets and Parameter Settings
To evaluate the algorithm, we mainly used four commonly used datasets: the Famous dataset [13], which contains 22 well-known models in geometric processing; MeshSegBenchmark (MeshSeg) [64] is a 3D mesh segmentation benchmark consisting of 380 meshes from 19 object categories. Since the meshes within each category share similar characteristics, we randomly selected 60 meshes for our experiments. We also use models from large datasets, such as ABC [65] and Thingi10K [66]. The ABC dataset contains approximately 1 million CAD models, while Thingi10K includes 10,000 3D printable meshes. Using all available models for evaluation would result in prohibitively high computational cost. To balance efficiency and representativeness, we follow the experimental setup of Points2Surf [13] and select a subset of 100 meshes from each dataset for our experiments. In addition, we also used real scanned data [13] to demonstrate the effectiveness of SE-RBFNet.
We set the parameters of SE-RBFNet as follows: batch size $B_{s}=10000$, max epoch $M^{e}=2000$, learning rate $lr=0.01$, maximum depth of the octree $l_{max}=10$, the two thresholds for filtering training points: $\tau_{1}=0.9$ and $\tau_{2}=1.1$, $\tau^{l_{1}}=10$ and $k^{l_{1}}=50$ are used to determine whether the number of basis functions has stabilize, while $\tau^{l_{2}}=0.5$ and $k^{l_{2}}=10$ are used to judge the stability of the $\mathcal{L}_{2}$ loss. $\tau_{m}=0.0005$ controls the sparsity optimization process, and $\tau_{d}=0.01$ is used as the deletion threshold for ERBFs. The initial training layers $l_{s}$ is set by default to the third-to-last layer. However, it must also satisfy the following condition: $\hat{S}^{l_{s}}_{\text{in}}>100$, otherwise, $l_{s}$ is incremented by one, i.e., $l_{s}=l_{s}+1$. This constraint ensures that the number of inner grid points in the $l_{s}$ layer must be greater than 100 to avoid having too few initial ERBFs. Experiments have found that $l_{s}$ is generally 6 or 7. The specific use of all parameters is shown in Algorithm 3. For the comparison method SparseRBF, we adopt the default parameter settings as reported in its original paper [1].

### v4

@@ 3.4.3 Adaptive Basis Function Addition
We introduce an adaptive basis function addition mechanism that automatically adds suitable basis functions to areas with large errors during the optimization of the loss function. Before adding new basis functions, two prerequisites must be satisfied: (1) the condition in Eq. (12) has been fulfilled, meaning the $\mathcal{L}_{2}$ loss has stabilized; (2) the number of effective basis functions remains stable over a predefined number of optimization steps, formulated as:
@@
where $b_{f}^{(e)}$ indicates the number of effective basis functions at epoch $e$, and is computed as: $b_{f}^{(e)}=\sum_{j=1}^{M}\mathbb{I}\left(\left|w_{j}\right|\geq\tau_{d}\right)$, $\mathbb{I}(\cdot)$ is the indicator function, which returns 1 if the condition is true and 0 otherwise. The threshold $\tau_{d}$ is used to determine whether a basis function is effective. $k^{l_{1}}$ and $\tau^{l_{1}}$ represent the number of iterations used to determine the stability of the number of basis functions and the corresponding threshold, respectively. Once the condition of Eq. (19) is triggered, the basis function addition process is automatically performed. After adding the new basis functions, the $\mathcal{L}_{1}$ regularization term is temporarily deactivated and the network switches to pure $\mathcal{L}_{2}$ optimization to fully train the newly added basis functions. The $\mathcal{L}_{1}$ regularization term is reactivated for sparse optimization when Eq. (12) and Eq. (13) are satisfied again. The process repeats as follows:
@@
    When Eq. (12) and Eq. (13) are satisfied, $\mathcal{L}_{1}$ regularization is reactivated for sparse optimization.
@@
1:Identify high-error points: $\mathcal{I}_{add}=\{i\mid|E_{i}|>\frac{\tau_{m}}{2}\}$.
@@
11:  return $\emptyset,\emptyset,\emptyset,\emptyset$
@@
The process alternates cyclically until the predefined number of training epochs $T^{e}$ is reached. The detailed process for adding basis functions is given in Algorithm 2. First, calculate the error vector for each training point in $V^{ls}$, and select the subset indexed by $id_{all}$, which is computed using Eq. (10):
$\bm{A}^{a}=\mathbf{0},$
where $\mathcal{I}_{ext}$ denotes the indices of the extreme error points. The length of $\mathcal{I}_{ext}$ is $k^{a}$, which indicates the number of newly added basis functions. $\bm{C}^{a}\in\mathbb{R}^{k^{a}\times 3}$ represents the centers of the newly added basis functions. $\bm{W}^{a}\in\mathbb{R}^{k^{a}\times 1}$ represents the corresponding weights. $\bm{d}^{a}\in\mathbb{R}^{k^{a}}$ represents the axis lengths of the newly added basis functions. Since the axis lengths are initialized to be identical along all three directions, each scalar element in $\bm{d}^{a}$ is expanded into a $3\times 3$ diagonal matrix to construct $\bm{D}^{a}\in\mathbb{R}^{k^{a}\times 3\times 3}$. $\bm{A}^{a}\in\mathbb{R}^{k^{a}\times 3}$ denotes the rotation angles, which are initialized to zero. $\bar{d}_{i}$ refers to the $i$-th element of $\bm{\bar{d}}$, which is defined as the minimum distance from the center $\bm{c}^{a}(i)$ to the sampling points $P$, $\bar{d}_{i}=\min\left\|\bm{c}^{a}(i)-P\right\|_{2},i=1,2,\ldots,k^{a}$. $\operatorname{sign}()$ is the sign function:
$\operatorname{sign}(x)=\begin{cases}-1,&x<0\\ 0,&x=0\\ 1,&x>0\end{cases}.$
@@
During SE-RBFNet optimization, some basis functions have coefficients that become so small that their contribution to the overall output can be ignored. We introduce a threshold $\tau_{d}$ and remove the basis functions whose $w$ coefficients are less than $\tau_{d}$ at the specified iteration interval. That is:
$\bm{\widetilde{W}}$ $=\left\{w_{j}\in\bm{W}|\>|w_{j}|<\tau_{d}\right\},$ (24)
$\bm{\widetilde{C}}$ $=\left\{\bm{c}_{j}\in\bm{C}|\>|w_{j}|<\tau_{d}\right\},$
$\bm{\widetilde{D}}$ $=\left\{\bm{D}_{j}\in\bar{\bm{D}}|\>|w_{j}|<\tau_{d}\right\},$
$\bm{\widetilde{A}}$ $=\left\{\bm{a}_{j}\in\bm{A}|\>|w_{j}|<\tau_{d}\right\}.$
Deleting these basis functions allows the network to automatically adjust its complexity, increasing sparsity and computational efficiency. When training reaches $T^{e}$ epochs, the optimization focuses exclusively on the $\mathcal{L}_{2}$ loss term to further improve the precision of the SDF approximation, continuing until the maximum iteration count $M^{e}$ is reached. At this stage, no further basis function additions or deletions occur, and $\mathcal{L}_{1}$ regularization is no longer applied.
3.4.4 Additional Techniques for Training Acceleration
In practice, the size of the input set $G$ contains a very large number of points, and the problem of optimization $\mathcal{L}$ is a high-dimensional non-convex nonlinear problem. To reduce computational burden and increase convergence speed, we primarily employ the following techniques for accelerating the training process.
@@ 3.4.4 Additional Techniques for Training Acceleration
$\bm{F}_{ij}=\begin{cases}\hat{\Phi}\left(\bm{v}_{i},\Theta_{j}\right),&\text{ if }\left\|\bm{c}_{j}-\bm{v}_{i}\right\|_{2}<\sqrt{\frac{-\ln(\varepsilon)}{\lambda_{j}}}\\ 0,&\text{ otherwise. }\end{cases}$ (26)
Algorithm 3 SE-RBFNet Sparse Optimization Algorithm
Input: Sampling points $P$ on an implicit surface, batch size $B_{s}=10000$, $M^{e}=2000$, $T^{e}=1600$, initial learning rate $lr=0.01$, maximum depth of the octree $l_{max}=10$, $l_{s}$ is set by default to the third-to-last layer, $\gamma=10^{-3}$, $\varepsilon=10^{-7}$, $\tau_{1}=0.9$, $\tau_{2}=1.1$, $\tau_{m}=0.02$, $\tau_{d}=0.01$, $\tau^{l_{1}}=5$, $\tau^{l_{2}}=0.5$, $k^{l_{1}}=50$, $k^{l_{2}}=10$. 1:Construct an octree from the sampling points $P$, and compute the SDF at each octree grid point $G$, which is then normalized according to Eq. (6) to obtain $\hat{S}$ as the target for sparse optimization; 2:According to the $l_{s}$, obtain the initial training points $V^{l_{s}}$ and corresponding real labels $T^{ls}$, and then calculate the parameters of the initial ERBF using Algorithm 1. 3:Initialize SE-RBFNet parameters: $\bm{\widetilde{\Theta}}\leftarrow\{\bm{C}\textrm{,}\bar{\bm{D}}\textrm{,}\bm{A},\bm{W}\}$. 4:Set $l_{1}\_{optim}$=False, add_pointadd\_point=False, $e_{p}=1$; 5:while $e_{p}<M^{e}$ do 6:  $e_{p}\leftarrow e_{p}+1$, 7:  if add_pointadd\_point and $l_{s}<l_{max}$ then 8:   The grid points from the $(l_{s}+1)$-th layer are selected and appended to the training set, followed by updating: $V^{l_{s}}\leftarrow V^{l_{s}+1}$, $T^{ls}\leftarrow T^{ls+1}$ and $l_{s}\leftarrow l_{s}+1$. 9:   Set add_pointadd\_point = False. 10:  end if 11:  Shuffle training points and corresponding labels. 12:  for $i=0$ to $N$ with step $B_{s}$ do 13:   Sample batch $V^{l_{s}}_{b}$ and $T^{l_{s}}_{b}$ 14:   Use Eq. (5) to calculate the network forward result: $\bm{O}\leftarrow\text{SE-RBFNet}(V^{l_{s}}_{b})$ 15:   Calculate $id_{all}$ according to Eq. (10) 16:   Compute $\mathcal{L}_{2}$ and $\mathcal{L}_{1}$ loss using Eq. (9) 17:   Explicitly compute the gradient $\nabla{\mathcal{L}(\bm{\widetilde{\Theta}})}$ using the formulas in A. 18:   if $l_{1}\_{optim}$=True and Eq. (13) is True then 19:     Use Eq. (14) to compute adaptive $\alpha,\beta$ 20:     Update loss: $\mathcal{L}\leftarrow\alpha\mathcal{L}_{2}+\beta\mathcal{L}_{1}$ 21:   else 22:     Update loss: $\mathcal{L}\leftarrow\mathcal{L}_{2}$ 23:   end if 24:   Update network parameters $\bm{\widetilde{\Theta}}$ via Adam [28]. 25:  end for 26:  if $e_{p}\>\%\>k^{l_{2}}=0$ and $l_{1}\_{optim}$=True and $e_{p}<T^{e}$ then 27:   Delete invalid basis functions using Eq. (24). 28:  end if 29:  if $l_{1}\_{optim}$=False and Eq. (12) is True then 30:   Update $l_{1}\_{optim}$=True. 31:  end if 32:  if $l_{1}\_{optim}$=True and Eq. (19) is True then 33:   Add new ERBF according to Eq. 21. 34:   Update $l_{1}\_{optim}$=False. 35:  end if 36:  if $e_{p}=T^{e}$ then 37:   During the iterations from $T^{p}$ to $M^{p}$, steps 26 to 35 are skipped, and only $\mathcal{L}_{2}$ optimization is performed. $l_{1}\_{optim}$=False, $lr=10^{-3}$ and adjusted using a cosine annealing schedule, with a minimum value of $10^{-5}$. 38:  end if 39:  if $e_{p}$ in $[400,800]$ then 40:   $add\_points=\text{True}$ 41:  end if 42:end while Output: Network parameters $\bm{\widetilde{\Theta}}$, predicted SDF $\bm{O}$, a surface via the marching cubes algorithm [34].
@@
The detailed derivations of the gradients can be found in A. As shown in the A, when the value of $\bm{F}$ at a certain location is close to zero, the corresponding gradient also approaches zero. Consequently, to improve the calculation efficiency, we take advantage of the nearest-neighbor strategy as described in Eq. (25), only the gradient values of the points near the center points of the basis functions are calculated.
Finally, due to the inherent independence of each ERBF in the network, we leverage CUDA to parallelize the computation of the value of the loss function and the gradient with respect to each parameter in $\Theta$. This independence allows each ERBF to be computed independently and concurrently, leading to a significant reduction in computational time.
@@ 4 Experiment and analysis
In this section, we evaluate the accuracy, sparsity, and efficiency of SE-RBFNet. It is important to emphasize that SE-RBFNet is designed to approximate the SDF of a general implicit surface, rather than directly reconstruct surfaces from raw point clouds. Therefore, we do not perform direct comparisons with full surface reconstruction methods. For experimental demonstration, we conducted two complementary experiments: one using ground-truth SDFs from triangle meshes, and the other using predicted SDFs from neural implicit methods (or any other SDF-generating methods).
First, to assess sparsity and parameter efficiency, we compare SE-RBFNet with SparseRBF [30], which also employs sparse RBFs for implicit surface representation. In the original SparseRBF framework, the inputs consist of surface points (with zero SDF), offset points, and center points, where the SDF values are approximated from Voronoi diagram poles [1] rather than computed exactly. To ensure a fair comparison, we instead provide both methods with identical input data: area-weighted uniformly sampled points on the mesh surface (SDF = 0) and octree grid points, whose SDF values are defined as the directional point-to-mesh distances and computed directly from the triangle mesh geometry using the GPU-accelerated Kaolin API [22]. The octree construction follows the implementation of [31]. This setup eliminates potential inaccuracies introduced by approximate SDF estimation in SparseRBF and ensures that the comparison focuses solely on the sparsity and approximation accuracy of the two methods.
Second, to demonstrate the generality of SE-RBFNet, we evaluate its ability to sparsely approximate arbitrary SDF data without emphasizing the source and the accuracy of the SDF computation. In this scenario, we start from sampled points on the implicit surface to construct the corresponding octree and obtain SDF values on the octree grid points using neural implicit methods or arbitrary SDF-generation methods. SE-RBFNet further sparsifies the representation by approximating the input SDFs using significantly fewer parameters, while preserving high accuracy of the surface from the approximated SDF by SE-RBFNet. This highlights the compatibility of SE-RBFNet with existing SDF generation methods and its advantage in reducing storage and transmission costs.
Together, these two parts of the experiments allow us to evaluate both the sparsity and parameter efficiency of SE-RBFNet, as well as its generality across different SDF sources. All experiments were performed on Ubuntu 18.04 with an Intel Core i9-7940X CPU @ 3.1GHz and an NVIDIA GeForce RTX 2080 GPU. For reproducibility, the executable used in our experiments is available at https://github.com/lianbobo/SE-RBFNet.git
@@ 4.1 Datasets and Parameter Settings
To evaluate the algorithm, we mainly used four commonly used datasets: the Famous dataset [15], MeshSegBenchmark (MeshSeg) [7], ABC [29], and Thingi10K [62]. The Famous dataset, which contains 22 well-known models in geometric processing, and MeshSeg, a 3D mesh segmentation benchmark consisting of 380 meshes from 19 object categories. Since the meshes within each category share similar characteristics, we randomly selected 60 meshes for our experiments. The ABC dataset contains approximately 1 million CAD models, while Thingi10K includes 10,000 3D printable meshes. Using all available models for evaluation would result in a prohibitively high computational cost. To balance efficiency and representativeness, we follow the experimental setup of Points2Surf [15] and select a subset of 100 meshes from each dataset for our experiments. In addition, we also used several real scanned data [15] to demonstrate the effectiveness of SE-RBFNet.
We set the parameters of SE-RBFNet as follows: batch size $B_{s}=10000$, max epoch $M^{e}=2000$, learning rate $lr=0.01$, maximum depth of the octree $l_{max}=10$, the two thresholds for filtering training points: $\tau_{1}=0.9$ and $\tau_{2}=1.1$, $\tau^{l_{1}}=5$ and $k^{l_{1}}=50$ are used to determine whether the number of basis functions has stabilized, while $\tau^{l_{2}}=0.5$ and $k^{l_{2}}=10$ are used to judge the stability of the $\mathcal{L}_{2}$ loss. $\tau_{m}=0.02$ controls the sparsity optimization process, and $\tau_{d}=0.01$ is used as the deletion threshold for ERBFs. The initial training layer $l_{s}$ is set by default to the third-to-last layer. However, it must also satisfy the following condition: $\hat{S}^{l_{s}}_{\text{in}}>100$, otherwise, $l_{s}$ is incremented by one, i.e., $l_{s}=l_{s}+1$. This constraint ensures that the number of inner grid points in the $l_{s}$ layer must be greater than 100 to avoid having too few initial ERBFs. Experiments have found that $l_{s}$ is generally 6 or 7. The specific use of all parameters is shown in Algorithm 3. All comparison methods are used with their default parameter settings.